# Robust Deterministic Policy Gradient for Disturbance Attenuation and Its Application to Quadrotor Control*

### v1

@@ III-A $H_{\infty}$ Control
Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@ IV-A Robust deterministic policy gradient
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{R^{n}}$
@@
${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
@@
${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}=E\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim{\rho^{{\pi_{\theta}},{\mu_{\phi}}}}}\right]$
@@ V-A Quadrotor dynamics model
where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame. Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame. Moreover, $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows:

### v3

@@ III-A $H_{\infty}$ Control
where $x_{k}\in\mathbb{R}^{n}$ is the state, $u_{k}\in\mathbb{R}^{m}$ is the control input, $w_{k}\in\mathbb{R}^{p}$ is the disturbance, and $v_{k}\in\mathbb{R}^{q}$ is the process noise at time step $k$. Using the state-feedback controller $u=\pi(x)$, the system can be reduced to the autonomous closed-loop system
@@ IV-A Robust deterministic policy gradient
$u=\pi_{\theta}(x)\in{\mathbb{R}}^{m},\quad w=\mu_{\phi}(x)\in{\mathbb{R}}^{p},\quad x\in{\mathbb{R}^{n}}$
@@
${\nabla_{\theta}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\theta}}{Q^{\pi,{\mu_{\phi}}}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\pi={\pi_{\theta}}}}}\right|s\sim\rho}\right]$
@@
${\nabla_{\phi}}{J^{{\pi_{\theta}},{\mu_{\phi}}}}={\mathbb{E}}\left[{\left.{{{\left.{{\nabla_{\phi}}{Q^{{\pi_{\theta}},\mu}}(s,{\pi_{\theta}},{\mu_{\phi}})}\right|}_{\mu={\mu_{\phi}}}}}\right|s\sim\rho}\right]$
@@ V-A Quadrotor dynamics model
where $m$ is the mass of the quadrotor, $g$ is the gravity acceleration, $T_{s}$ is the sampling time, $w_{k}=[w_{k,x},w_{k,y},w_{k,z}]^{T}$ means an external disturbance applied to the quadrotor, $\mathbf{a}_{k}=[a_{k,x},a_{k,y},a_{k,z}]^{T}$ denotes linear acceleration, $\mathbf{v}_{k}=[v_{k,x},v_{k,y},v_{k,z}]^{T}$ represents linear velocity, $\mathbf{p}_{k}=[p_{k,x},p_{k,y},p_{k,z}]^{T}$ describes the position of the quadrotor in the inertial frame, $\Delta\omega_{k}=[\Delta\omega_{k,\phi},\Delta\omega_{k,\theta},\Delta\omega_{k,\psi}]^{T}$ is angular acceleration, $\omega_{k}=[\omega_{k,\phi},\omega_{k,\theta},\omega_{k,\psi}]^{T}$ is the angular velocity, and $\rho_{k}=[\phi_{k},\theta_{k},\psi_{k}]^{T}$ represents the roll, pitch, yaw of the quadrotor in the body frame. Moreover, $\mathbf{R}$ is the rotation matrix from the body frame to the inertial frame, and $\mathbf{T}_{B}$ is the total thrust of the quadrotor along the $z$ axis in the body frame and defined as follows: